# Hermes: Memory-Efficient Pipeline Inference for Large Models on Edge Devices

### v1

@@ I Introduction
Due to the explosive growth in the size of large models, deploying them at the edge faces critical memory challenges [16]. Specifically, current edge devices offer only a limited amount of memory capacity, ranging from a few tens of megabytes to a few gigabytes. For example, NVIDIA Jetson Nano has 4 GB of memory and Raspberry Pi 4 Model B has up to 8 GB of memory. In contrast, large models’ parameters have experienced exponential growth, reaching sizes in the hundreds of billions. For instance, the GPT-3 [17] model has 175 billion trainable parameters, while the recently developed GPT-4 model exceeds the trillion parameter mark. Consequently, the memory usage of these large models can easily reach tens to hundreds of gigabytes, far surpassing the memory capacity of typical edge devices.
Existing works to address the memory challenges of large model inference on edge devices can be classified into two categories. The first attempts to optimize the model structure to reduce the computational load through techniques including model pruning [18, 19], model compression [20, 21], model quantization [22] and adaptive inference [23, 24]. Although these approaches significantly diminish the number of required computational operations, they often result in reduced model accuracy. Moreover, these approaches are generally tailored for specific models, thus limiting their applicability on a broader scale. The second optimizes the memory usage during model inference by model swapping between different storage media [25, 26, 27]. This method initially divides the model into separate shards and selectively preloads certain shards from disk into a buffer or the memory as needed for the inference process. Even though memory swapping methods can reduce memory overhead, they can inadvertently prolong inference latency due to the increased frequency of I/O operations between varying storage media.
@@
To resolve two issues mentioned above, we develop PipeLoad, a memory-efficient pipeline execution mechanism to streamline the loading and inference process with per-layer granularity. Building upon this mechanism, we introduce Hermes, an innovative framework for large model inference on edge devices. This framework integrates three main components. Firstly, the Layer Profiler evaluates the performance and memory utilization of each individual model layer for a given transformer model. Next, the Pipeline Planner utilizes the profiling data from Layer Profiler to devise a PipeLoad execution schedule within different memory constraints. Lastly, the Execution Engine determines the execution strategy from the schedule based on the current memory constraints of the edge device and executes PipeLoad inference.
@@ II-A Transformer Model Structure
The architectural makeup of transformer models is the basis for developing pipeline inference strategies on edge devices. A typical transformer model comprises embedding, encoder, decoder, pooling, and additional specialized layers. Based on their architectural configurations, they can be classified into three primary categories: encoder-decoder, encoder-only and decoder-only models. Encoder-decoder models, such as BART [30] and T5 [31] integrate both encoder and decoder layers. BART is architected for complex sequence-to-sequence tasks, and T5 generalizes this capability with a comprehensive text-to-text methodology applicable to a wide array of NLP challenges. Encoder-only models like BERT employ a series of encoder layers to interpret input data, ideally suited for tasks that do not involve generating sequences. Meanwhile, Vision Transformer (ViT) presents an innovative adaptation of encoder-only architecture to analyze sequences of image patches, eschewing the conventional decoder layout found in text-centric models. Conversely, decoder-only models such as GPT rely exclusively on decoder layers, focusing on the generation of new content by drawing on recognizable patterns.

### v2

@@ I Introduction
Due to the explosive growth in the size of large models, deploying them at the edge faces critical memory challenges [16]. Specifically, current edge devices offer only a limited amount of memory capacity, ranging from a few tens of megabytes to a few gigabytes. For example, NVIDIA Jetson Nano has 4 GB of memory and Raspberry Pi 4 Model B has up to 8 GB of memory. In contrast, large models’ parameters have experienced exponential growth, reaching sizes in the hundreds of billions. For instance, the GPT-3 [17] model has 175 billion trainable parameters, while the recently developed GPT-4 [18] model exceeds the trillion parameter mark. Consequently, the memory usage of these large models can easily reach tens to hundreds of gigabytes, far surpassing the memory capacity of typical edge devices.
Existing works to address the memory challenges of large model inference on edge devices can be classified into two categories. The first attempts to optimize the model structure to reduce the computational load through techniques including model pruning [19, 20, 21], model compression [22, 23], model quantization [24] and adaptive inference [25, 26]. Although these approaches significantly diminish the number of required computational operations, they often result in reduced model accuracy. Moreover, these approaches are generally tailored for specific models, thus limiting their applicability on a broader scale. The second optimizes the memory usage during model inference by model swapping between different storage media [27, 28, 29]. This method initially divides the model into separate shards and selectively preloads certain shards from disk into a buffer or the memory as needed for the inference process. Even though memory swapping methods can reduce memory overhead, they can inadvertently prolong inference latency due to the increased frequency of I/O operations between varying storage media.
@@
To resolve two issues mentioned above, we develop PipeLoad, a memory-efficient pipeline execution mechanism to streamline the loading and inference process with per-layer granularity. This mechanism incorporates dynamic management of memory with timely destruction of model weights that have been inferred, significantly reducing the memory usage of model inference. And by engaging parallel model loading, PipeLoad overlaps multiple inference time within a single loading interval to minimise the pipeline stalls, consequently decreasing inference latency.
Building upon this mechanism, we introduce Hermes, an innovative framework for large model inference on edge devices. This framework integrates three main components. Firstly, the Layer Profiler evaluates the performance and memory utilization of each individual model layer for a given transformer model. Next, the Pipeline Planner utilizes the profiling data from Layer Profiler to devise a PipeLoad execution schedule within different memory constraints. Lastly, the Execution Engine determines the execution strategy from the schedule based on the current memory constraints of the edge device and executes PipeLoad inference.
@@ II-A Transformer Model Structure
The architectural makeup of transformer models is the basis for developing pipeline inference strategies on edge devices. A typical transformer model comprises embedding, encoder, decoder, pooling, and additional specialized layers. Based on their architectural configurations, they can be classified into three primary categories: encoder-decoder, encoder-only and decoder-only models. Encoder-decoder models, such as BART [32] and T5 [33] integrate both encoder and decoder layers. BART is architected for complex sequence-to-sequence tasks, and T5 generalizes this capability with a comprehensive text-to-text methodology applicable to a wide array of NLP challenges. Encoder-only models like BERT [34] employ a series of encoder layers to interpret input data, ideally suited for tasks that do not involve generating sequences. Meanwhile, Vision Transformer (ViT) presents an innovative adaptation of encoder-only architecture to analyze sequences of image patches, eschewing the conventional decoder layout found in text-centric models. Conversely, decoder-only models such as GPT rely exclusively on decoder layers, focusing on the generation of new content by drawing on recognizable patterns.